find dalta PH of 400 ml , and 1.17 molar NaOH ,when the same amount of ,100 ml of 0.05 Molar NaOH added


1
Expert's answer
2022-02-01T03:01:01-0500

Initial pH"=14-(-log\>1.17)"

"=14.068"


Moles of OH"^-" added

"=\frac{100}{1000}×0.05=0.005"


Moles in 400ml of 1.17 molar NaOH

"=\frac{400}{1000}×1.17=0.468"


Total moles"=0.468+0.005"

"=0.473"


Total volume"=500ml"


Concentration "=\frac{1000}{500}×0.473"

"=0.946"


Final pH"=14-(-log\>0.946)"

"=13.976"




Change in pH"=14.068-13.976"

"=0.092"